OpenStudy (anonymous):

Medal for help :) 1/4a = 3 A. a = 0.75 B. a = 3 C. a = 12****My answer D. a = 16

OpenStudy (anonymous):

@ARMYRANGERS @Awesome781 @dan815

OpenStudy (anonymous):

1/4a = 3 1/4a = a/4 a/4 = 3 a=3(4) a=12

OpenStudy (anonymous):

so I was correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok, thank you